Prev Next

SW&IT Meeting Minutes 2022-04-18

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • Career Conversations; due by June 30
  • asm/rigid f/w instructions (step 45) update

Team reports

MB

  • Last week
    • xxx
  • This week
    • xxx

YZ

  • Previous:
    • Some LBC work and discussions with Olga (IT7847 and FPIA).
    • Other issuetrak investigation (mods1data freeze, TCS shared memory problems).
    • Worked with Guido and Xianyu updating DX hexapod firmware to adsec.
    • Worked with Alessandro and Olga on LBC bias level adjustment.
    • Archive meeting on Thursday.
    • TMS meeting and working on a new software request.
  • Next:
    • Continue LBC work with Olga.
    • Continue some issuetrak work (mods1data freeze and TCS shared memory).
    • Release a special version of TMS software (and maybe some test?).
    • Work on some MODS software issue if have time.

DE

  • Last week:
    • Archive meeting
      • Agreed on timing for archive-in work: Tuesday 19
    • Veeam license get and install
  • This week:
    • Prep for newdata swap
    • Actually do newdata swap
    • M1 Mac Mini should be coming soon, set up for development
    • Ketiv housekeeping

GE

  • Last week:
    • Spent a day and a quarter on medical appointments.
    • Put together a working group on upgrading the LUCI GUIs.
    • Continued poking around my TCS VM to upgrade from 2018A to the latest version.
    • Worked on IT 8573 with Matthieu.
    • Worked on ITs related to the Observer Scheduler GUI (8571, 8529).
  • This week:
    • Finally finish upgrading my TCS VM to the latest version.
    • Work on Observer Scheduler GUI (8571, 8529).

BM

  • Last week:
    • lbtidx bcu communication issues resolved thanks everyone
    • interrupted offsets (met with JCG/MB Monday, AO-SW mtg Friday)
      • new questions for Alfio (masterdiagnostic frame downloading during open loop, disturb enable/disable during wfsarb open loop, camera lens tracking recovery during safeskips)
    • IT 8602 - new test plan for CopleyCtrl fix (report errors back to wfsarb instead of timing out when movements are out of range)
    • argos: updated POC for LCH scripts--will discuss where these scripts should reside/Github w/Matthieu
  • This week:

PSG

  • Last week
    • NOMIC fixes and testing.
    • ARGOS final testing.
    • ARGOS deployment.
    • New version of OSS was activated by TO.
    • Update wiki with new OSS information.
  • This week
    • Go over ARGOS deployment - the process was not smooth.
    • Start examining AZCAM code base.
    • Further NOMIC testing - there may still be issues.
Topic revision: r9 - 18 Apr 2022, YangZhang
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback